OpenStudy (anonymous):

Explain how to make 200 mL of 2 M sodiun hydroxide (NaOH)

OpenStudy (anonymous):

Do you that C = n/V?

OpenStudy (anonymous):

i need to know how to do it

OpenStudy (anonymous):

But do you know that equation?

OpenStudy (anonymous):

(Its important in solving this problem)

OpenStudy (anonymous):

ok i understand that a mole is 6.02x10^23 of anything

OpenStudy (anonymous):

... Concentration ( mol / L) = n ( number of moles) / Volume ( L) You know both the concentration, and volume of the solution you want to make 2M is the equivalent to 2 mols per L And the volume you have is in mL, so you have to change it to L Using C = n/v, you can find the number of moles, and since you have the compound, you can find its' molar mass Knowing the fact that mass of a compound is the number of moles times molar mass. You can find the amount of sodium hydroxide (NaOH) needed to create the solution

OpenStudy (anonymous):

so i can use the equation C=n/v

OpenStudy (anonymous):

Yeah.. but you have to converl mL to L

OpenStudy (anonymous):

And rearrange for n, moles..

OpenStudy (anonymous):

ok give me one and ill try to do do beacuse im still not unstanding

OpenStudy (anonymous):

one what

OpenStudy (anonymous):

example like a gave before

OpenStudy (anonymous):

V = 200mL or 0.2L (divide by 100) C = 2M = 2mol/L C = n/V => n = CV, then n = (2mol/L)(0.2L), Litres cancel and you have mol => n = 0.4 mols, molar mass of NaOH = 22.99 g/mol + 16.00 g/mol + 1.01 g/mol To find out much mass you have to dissolve into 200mL of water to make a 2 mol/L solution of NaOH, you multiply 40g/mol * 0.4 mol to get 16 g

OpenStudy (anonymous):

how u get 0.2

OpenStudy (anonymous):

V = 200mL or 0.2L (divide by 1000)

OpenStudy (anonymous):

y u use 100

OpenStudy (anonymous):

type

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

1000 mL = 1 L...

OpenStudy (anonymous):

so i divide it by 1000

OpenStudy (anonymous):

Yes.. 200mL / 1000 gives you a number in L

OpenStudy (anonymous):

im sorry is there is there a way i can email u and u can explain it like that because there is no way i can get the full undwrstanding if im trying focus on this and thaat

OpenStudy (anonymous):

Id really prefer if we'd do it here

OpenStudy (anonymous):

im tryinh yo understand

OpenStudy (anonymous):

The thing is you want to make a solution that has a concetration of 0.2 moles per litre. You want do dissolve a certain mass of NaOH into 200mL of water to make that..

OpenStudy (anonymous):

ok i understand that

OpenStudy (anonymous):

next

OpenStudy (anonymous):

In order to find the mass of NaOH you want to dissolve, you have to find the number of moles you need. Using C = n/V, since you have the both the concentration and volume of the solution you want to make, you can find the number of moles that need to be dissolved

OpenStudy (anonymous):

ok give me a example like make up one

OpenStudy (anonymous):

Molecular weight of NaOH=40 g. Molarity = No of moles/litre If you want to preperae 2M,200 ml 2*40*0.2= 16 g Take 16g NaOH and dissolve in water in a standard flask upto 200ml.

OpenStudy (anonymous):

how iw u get 0.2

OpenStudy (anonymous):

200 ml=0.2 litre